# Determination of the stably free cancellation property for orders

### v2

@@ Abstract.
Let $K$ be a number field, let $A$ be a finite-dimensional semisimple $K$-algebra, and let $\Lambda$ be an $\mathcal{O}_{K}$-order in $A$. We give practical algorithms that determine whether or not $\Lambda$ has stably free cancellation (SFC). As an application, we determine all finite groups $G$ of order at most $383$ such that the integral group ring $\mathbb{Z}[G]$ has SFC.
@@ 1. Introduction
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$. Let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $\mathcal{O}$-order in $A$. For example, if $G$ is a finite group then the group ring $\mathcal{O}[G]$ is an $\mathcal{O}$-order in the group algebra $K[G]$. For $k\geq 0$, let $\Lambda^{(k)}$ denote the (left) $\Lambda$-module that is the direct sum of $k$ copies of $\Lambda$. The order $\Lambda$ is said to have stably free cancellation (SFC) if every finitely generated stably free $\Lambda$-module is in fact free, that is, if $X$ is a finitely generated $\Lambda$-module such that $X\oplus\Lambda^{(n)}\cong\Lambda^{(n+m)}$ for some $m,n\geq 0$, then $X\cong\Lambda^{(m)}$.
@@
One can also consider locally free cancellation (LFC), the definition of which we defer to §3.1. If $\Lambda$ has LFC then it also has SFC. Note, however, that there exist orders that have SFC but do not have LFC (see [Sme15] or [SV19]). We say that $A$ satisfies the Eichler condition if no simple component of $A$ is a totally definite quaternion algebra (see §3.2). Jacobinski [Jac68] showed that if $A$ satisfies the Eichler condition then $\Lambda$ has LFC and thus SFC. Smertnig–Voight [SV19] showed that in the case that $A$ is a totally definite quaternion algebra, the question of whether $\Lambda$ has SFC or LFC can be decided algorithmically. They then used this to classify all orders in totally definite quaternion algebras that have SFC, and of these, all those that have LFC. This classification builds on work of Vignéras [Vig76], Hallouin–Maire [HM06] and Smertnig [Sme15].
@@
For notational convenience, we henceforth abuse language by referring to direct sums of orders and algebras rather than direct products. We can always decompose $A$ into a direct sum of $K$-algebras $A=A_{1}\oplus A_{2}\oplus\cdots\oplus A_{k}$, where $A_{1}$ satisfies the Eichler condition and $A_{2},\ldots,A_{k}$ are all totally definite quaternion algebras. If $\Lambda$ decomposes as $\Lambda=\Lambda_{1}\oplus\Lambda_{2}\oplus\cdots\oplus\Lambda_{k}$ such that each $\Lambda_{i}$ is contained in $A_{i}$, then the question of whether $\Lambda$ has SFC or LFC can be decided using the results discussed in the previous paragraph, since both properties respect direct sums. However, in many cases of interest, $\Lambda$ does not exhibit such a decomposition.
@@
In this article, we show that the question of whether an arbitrary order $\Lambda$ has SFC or not can in principle always be decided algorithmically. In fact, we present three algorithms, all of which are practical in certain situations. The first of these, Algorithm 8.9, is only practical when $A$ is of ‘small’ dimension. The second, Algorithm 9.1, cannot prove that $\Lambda$ has SFC, but can be used to show that it fails SFC, and in this situation is often much faster than Algorithm 8.9. The third, Algorithm 10.3, uses fibre products to reduce the problem for the original order to that for an order in an algebra of smaller dimension as well as the computation of a certain subgroup of the unit group of a finite ring. These algorithms require several other algorithms of independent interest, including those that can be used to compute certain primary decompositions and unit groups of finite rings (see §11), and new freeness tests for certain modules over orders (see §12).
@@
Orders of particular interest are integral group rings $\mathbb{Z}[G]$ for $G$ a finite group. For instance, the property of whether or not $\mathbb{Z}[G]$ has SFC or LFC has applications in topology (see [Joh03], [Nic21b], [Nic24a], for example) and in the determination of whether a finite Galois extension of $\mathbb{Q}$ has a normal integral basis or not (see [Cou94], [Cou98]).
@@
Let $G$ be a finite group and let $\mathbb{H}$ denote the quaternion division algebra over $\mathbb{R}$. If no simple component of $\mathbb{R}[G]$ is isomorphic to $\mathbb{H}$, then $\mathbb{Q}[G]$ satisfies the Eichler condition and so $\mathbb{Z}[G]$ has LFC (and thus SFC) by the aforementioned result of Jacobinski. This condition on $G$ is equivalent to the requirement that $G$ has no quotient that is isomorphic to a binary polyhedral group, that is, a generalised quaternion group $Q_{4n}$ for $n\geq 2$, or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups. In §4, we review important work of Swan [Swa83], Chen [Che86], and Nicholson [Nic21a] on SFC and LFC for $\mathbb{Z}[G]$ in the case that $G$ does have a quotient isomorphic to a binary polyhedral group.
@@
Using Algorithm 10.3 with judicious choices of fibre products, we prove the following result on new cases of integral group rings with SFC (this is Theorem 13.1).
@@
This represents a significant advance in the state-of-the-art, since before the results of this article, such a classification was only known for groups $G$ with $|G|\leq 31$ (for example, it was not known whether $\mathbb{Z}[G_{(32,14)}]$ had SFC or not). In fact, Theorem 1.4 is a special case of Theorem 13.4, which gives a classification for $|G|\leq 1023$, but with $49$ explicitly listed gaps. In other words, of the $11759892$ groups $G$ with $|G|\leq 1023$, there are now only $49$ for which it is not known whether $\mathbb{Z}[G]$ has SFC or not.
@@
Acknowledgements
The authors are grateful to John Nicholson for helpful comments and discussions, including suggestions for which integral group rings to investigate initially. They would also like to thank Nigel Byott for several corrections and suggestions. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any author accepted manuscript version arising.
@@ 2. Preliminaries on lattices and orders
For further background on lattices and orders, we refer the reader to [CR81, Rei03]. Henceforth, all rings considered will be associative and unital.
@@ 2.1. Lattices and orders
Let $R$ be a Dedekind domain with field of fractions $K$. An $R$-lattice is a finitely generated torsion-free $R$-module, or equivalently, a finitely generated projective $R$-module. For any finite-dimensional $K$-vector space $V$, an $R$-lattice in $V$ is a finitely generated $R$-submodule $M$ in $V$. We define a $K$-vector subspace of $V$ by
$KM:=\{\alpha_{1}m_{1}+\alpha_{2}m_{2}+\cdots+\alpha_{r}m_{r}\mid r\in\mathbb{Z}_{\geq 0},\alpha_{i}\in K,m_{i}\in M\}$
and say that $M$ is a full $R$-lattice in $V$ if $KM=V$. We may identify $KM$ with $K\otimes_{R}M$.
Now let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $R$-order in $A$, that is, a subring of a $A$ that is an $R$-lattice such that $K\Lambda=A$. A $\Lambda$-lattice is a left $\Lambda$-module that is also an $R$-lattice.
2.2. Localisations and completions
@@
Let $\mathfrak{p}$ range over all maximal ideals of $R$. Let $R_{\mathfrak{p}}$ denote the localisation of $R$ at $\mathfrak{p}$, that is, the ring of $\mathfrak{p}$-integral elements in $K$. For an $R$-lattice $M$, define the localisation $M_{\mathfrak{p}}$ to be the $R_{\mathfrak{p}}$-lattice $R_{\mathfrak{p}}\otimes_{R}M$. Localising a $\Lambda$-lattice $X$ at $\mathfrak{p}$ yields a $\Lambda_{\mathfrak{p}}$-lattice $X_{\mathfrak{p}}$. Two $\Lambda$-lattices $X$ and $Y$ are said to be locally isomorphic, or in the same genus, if there is an isomorphism of $\Lambda_{\mathfrak{p}}$-lattices $X_{\mathfrak{p}}\cong Y_{\mathfrak{p}}$ for every $\mathfrak{p}$, in which case we write $X\vee Y$.
@@ 2.3. Locally free lattices and the principal genus
A $\Lambda$-lattice $X$ is said to be locally free of rank $n$ if $X\vee\Lambda^{(n)}$, where $\Lambda^{(n)}$ denotes the direct sum of $n$ copies of $\Lambda$. Note that every locally free $\Lambda$-lattice is projective over $\Lambda$ by [CR81, (8.19)].
Define the principal genus $g(\Lambda)$ to be the collection of locally free $\Lambda$-lattices of rank $1$. By [CR81, (31.7)], given $Y_{1},Y_{2},Y_{3}\in g(\Lambda)$, there exists $Y_{4}\in g(\Lambda)$ such that
$Y_{1}\oplus Y_{2}\cong Y_{3}\oplus Y_{4}.$ (2.1)
Moreover, by [CR81, (31.14)], given $Y\vee\Lambda^{(n)}$, there exists $X\in g(\Lambda)$ such that
$Y\cong\Lambda^{(n-1)}\oplus X.$ (2.2)
For $X\in g(\Lambda)$ we write $[X]_{\mathrm{iso}}$ for the isomorphism class of $X$ and define
@@ 2.4. Stable isomorphism
Let $P(\Lambda)$ denote the collection of finitely generated projective $\Lambda$-modules. For $X,Y\in P(\Lambda)$, we say that $X$ and $Y$ are stably isomorphic if
$X\oplus\Lambda^{(m)}\cong Y\oplus\Lambda^{(m)}\textrm{ for some }m\geq 0.$
A module $X\in P(\Lambda)$ is said to be stably free of rank $n$ if it is stably isomorphic to $\Lambda^{(n)}$. For $X\in P(\Lambda)$, let $[X]_{\mathrm{st}}$ denote the stable isomorphism class of $X$. Note that by [CR81, (6.15)] and the discussion in §2.2, stably isomorphic $\Lambda$-lattices are locally isomorphic. In particular, stably free lattices are locally free.
@@ 2.5. Locally free class groups
By (2.1), given $X_{1},X_{2}\in g(\Lambda)$ there exists $X_{3}\in g(\Lambda)$ such that $X_{1}\oplus X_{2}\cong\Lambda\oplus X_{3}$. It is straightforward to check that the stable isomorphism class $[X_{3}]_{\mathrm{st}}$ is uniquely determined by the classes $[X_{1}]_{\mathrm{st}}$ and $[X_{2}]_{\mathrm{st}}$. We define a binary operation on $\mathrm{Cl}(\Lambda)$ by
$[X_{1}]_{\mathrm{st}}+[X_{2}]_{\mathrm{st}}=[X_{3}]_{\mathrm{st}}\textrm{ whenever }X_{1}\oplus X_{2}\cong\Lambda\oplus X_{3},$
where $X_{1},X_{2},X_{3}\in g(\Lambda)$. This is a well-defined commutative and associative operation, with identity element $[\Lambda]_{\mathrm{st}}$. Moreover, inverses exist since by (2.1), given $X\in g(\Lambda)$ there exists $X^{\prime}\in g(\Lambda)$ such that $X\oplus X^{\prime}\cong\Lambda^{(2)}$. Therefore $\mathrm{Cl}(\Lambda)$ is an abelian group. Moreover, $\mathrm{Cl}(\Lambda)$ is finite since $\mathrm{LF}_{1}(\Lambda)$ is finite and the canonical map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $[X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is surjective.

### v3

@@ Abstract.
Let $K$ be a number field, let $A$ be a finite-dimensional semisimple $K$-algebra, and let $\Lambda$ be an $\mathcal{O}_{K}$-order in $A$. We give practical algorithms that determine whether $\Lambda$ has stably free cancellation (SFC). As an application, we determine all finite groups $G$ of order at most $383$ such that the integral group ring $\mathbb{Z}[G]$ has SFC.
@@ 1. Introduction
A ring $R$ is said to have stably free cancellation (SFC) if every finitely generated stably free (left) $R$-module is in fact free, that is, if $M$ is a finitely generated (left) $R$-module such that $M\oplus R^{\oplus n}\cong R^{\oplus(n+m)}$ for some $m,n\geq 0$, then $M\cong R^{\oplus m}$. In this article, we develop practical algorithms to determine whether a given order in a finite-dimensional semisimple algebra over a number field has SFC.
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$. Let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $\mathcal{O}$-order in $A$. For example, if $G$ is a finite group then the group ring $\mathcal{O}[G]$ is an $\mathcal{O}$-order in the group algebra $K[G]$. In addition to stably free cancellation for $\Lambda$, one can also consider locally free cancellation (LFC), whose definition we defer to §3.1. If $\Lambda$ has LFC then it also has SFC. Note, however, that there exist orders with SFC but without LFC (see [Sme15] or [SV19]).
@@
We say that $A$ satisfies the Eichler condition if no simple component of $A$ is a totally definite quaternion algebra (see §3.2). Jacobinski [Jac68] showed that if $A$ satisfies the Eichler condition then $\Lambda$ has LFC and thus SFC. Smertnig–Voight [SV19] showed that when $A$ is a totally definite quaternion algebra, the questions of whether $\Lambda$ has SFC or LFC can be decided algorithmically. They then used this to classify all orders in totally definite quaternion algebras that have SFC, and of these, all those that have LFC. This classification builds on work of Vignéras [Vig76], Hallouin–Maire [HM06] and Smertnig [Sme15].
@@
We can always decompose $A$ into a direct product of $K$-algebras $A=\prod_{i=1}^{k}A_{i}$, where $A_{1}$ satisfies the Eichler condition and $A_{2},\ldots,A_{k}$ are all totally definite quaternion algebras. If $\Lambda$ decomposes as a product of $\mathcal{O}$-orders $\Lambda=\prod_{i=1}^{k}\Lambda_{i}$ such that each $\Lambda_{i}$ is contained in $A_{i}$, then the questions of whether $\Lambda$ has SFC or LFC can be decided using the results discussed in the previous paragraph, since both properties respect direct products. However, in many cases of interest, $\Lambda$ does not exhibit such a decomposition. The desire to consider such cases motivates the following result, which is the main result of this article.
Theorem 1.1.
There exists an algorithm that, given an arbitrary $\mathcal{O}$-order $\Lambda$ in a finite-dimensional semisimple $K$-algebra, decides whether $\Lambda$ has SFC.
@@
In fact, we present three algorithms, all of which are practical in certain situations. The first of these, Algorithm 8.9, is only practical when $A$ is of ‘small’ dimension. This relies on Theorem 8.7 and works by computing a maximal order $\mathcal{M}$ containing $\Lambda$, a suitable two-sided ideal $\mathfrak{f}$ of $\mathcal{M}$ contained in $\Lambda$, and the image of $(\Lambda/\mathfrak{f})^{\times}$ in a certain ray class group, and then using this data to determine whether a finite number of explicit stably free ‘test lattices’ are in fact free over $\Lambda$. The main ideas employed utilize and build upon the results of [BB06] and [BHJ22]. The second, Algorithm 9.1, cannot prove that $\Lambda$ has SFC, but can be used to show that it fails SFC, and in this situation is often much faster than Algorithm 8.9. This works by sampling ‘random’ test lattices and checking if they fail to be free over $\Lambda$. The third, Algorithm 10.3, uses fiber products to reduce the problem for the original order to that for an order in an algebra of smaller dimension as well as the computation of a certain subgroup of the unit group of a finite ring. It uses Corollary 5.12, which itself crucially relies on results of Reiner–Ullom [RU74] (see Proposition 5.2) and Swan [Swa80] (see Theorem 5.10). These algorithms require several other algorithms of independent interest, including those that can be used to compute certain primary decompositions and unit groups of finite rings (see §11), and new freeness tests for certain modules over orders (see §12).
@@
Orders of particular interest are integral group rings $\mathbb{Z}[G]$ for $G$ a finite group. For instance, the property of whether $\mathbb{Z}[G]$ has SFC or LFC has applications in topology (see [Joh03], [Nic21b], [Nic24a], for example) and in the determination of whether a finite Galois extension of $\mathbb{Q}$ has a normal integral basis (see [Cou94], [Cou98]).
@@
Let $G$ be a finite group and let $\mathbb{H}$ denote the quaternion division algebra over $\mathbb{R}$. If no simple component of $\mathbb{R}[G]$ is isomorphic to $\mathbb{H}$, then $\mathbb{Q}[G]$ satisfies the Eichler condition and so $\mathbb{Z}[G]$ has LFC (and thus SFC) by the aforementioned result of Jacobinski. This condition on $G$ is equivalent to the requirement that $G$ has no quotient that is isomorphic to a binary polyhedral group, that is, a generalized quaternion group $Q_{4n}$ for $n\geq 2$, or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups. In §4, we review important work of Swan [Swa83], Chen [Che86], and Nicholson [Nic21a] on SFC and LFC for $\mathbb{Z}[G]$ in the case that $G$ does have a quotient isomorphic to a binary polyhedral group.
@@
Using Algorithm 10.3 with judicious choices of fiber products, we prove the following result on new cases of integral group rings with SFC (this is Theorem 13.1).
@@
This represents a significant advance in the state-of-the-art, since before the results of this article, such a classification was only known for groups $G$ with $|G|\leq 31$ (for example, it was not known whether $\mathbb{Z}[G_{(32,14)}]$ had SFC). In fact, Theorem 1.5 is a special case of Theorem 13.4, which gives a classification for $|G|\leq 1023$, but with $49$ explicitly listed gaps. In other words, of the $11759892$ groups $G$ with $|G|\leq 1023$, there are now only $49$ for which it is not known whether $\mathbb{Z}[G]$ has SFC.
@@
Organization
This article is organized as follows. In §2, we give preliminaries on lattices and orders. We review cancellation properties for orders in §3, and in §4, we specialize to the case of integral group rings. In §5, we give reduction criteria for SFC for orders using fiber products, and in §6 we consider applications to integral group rings. Next, in §7, we give criteria for freeness and stable freeness of lattices over orders. We develop the three algorithms discussed above for determining whether an order has SFC in §8, §9 and §10. In §11, we give algorithms to compute certain primary decompositions and unit groups of finite rings, and in §12 we give new freeness tests for certain modules over orders. Our main results on SFC for integral group rings, which include Theorem 1.5 as a special case, are proven in §13. Finally, in Appendix A, we discuss the implementation of the algorithms of this article, and give relevant data and timings.
Acknowledgments
The authors are grateful to John Nicholson for helpful comments and discussions, including suggestions for which integral group rings to investigate initially. They would also like to thank Nigel Byott for several corrections and suggestions, as well as an anonymous referee for suggesting a number of improvements. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any author accepted manuscript version arising.
@@ Acknowledgments
Notation and conventions
All rings are assumed to be associative and unital. All modules are assumed to be left modules unless otherwise stated. For notational convenience, we shall henceforth abuse notation by using the symbol $\oplus$ to denote direct products of orders and algebras. Denote the center of a ring $R$ by $Z(R)$. For a division algebra $D$ and a positive integer $k$, let $\mathrm{Mat}_{k}(D)$ denote the ring of all $k\times k$ matrices with entries in $D$. We call $\mathrm{Mat}_{k}(D)$ a full matrix algebra over $D$.
@@ 2. Preliminaries on lattices and orders
For further background on lattices and orders, we refer the reader to [CR81, Rei03].
@@ 2.1. Lattices and orders
Let $R$ be a Dedekind domain with field of fractions $K$. An $R$-lattice is a finitely generated torsion-free $R$-module, or equivalently, a finitely generated projective $R$-module. For any finite-dimensional $K$-vector space $V$, an $R$-lattice in $V$ is a finitely generated $R$-submodule $M$ of $V$. We say that $M$ is a full $R$-lattice in $V$ if $KM=V$. We may identify $K\otimes_{R}M$ with $KM$ via the map $k\otimes m\mapsto km$.
Now let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $R$-order in $A$, that is, a subring of $A$ that is an $R$-lattice such that $K\Lambda=A$. A $\Lambda$-lattice is a left $\Lambda$-module that is also an $R$-lattice.
2.2. Localizations and completions
@@
Let $\mathfrak{p}$ range over all maximal ideals of $R$. Let $R_{\mathfrak{p}}$ denote the localization of $R$ at $\mathfrak{p}$, that is, the ring of $\mathfrak{p}$-integral elements in $K$. For an $R$-lattice $M$, define the localization $M_{\mathfrak{p}}$ to be the $R_{\mathfrak{p}}$-lattice $R_{\mathfrak{p}}\otimes_{R}M$. Localising a $\Lambda$-lattice $X$ at $\mathfrak{p}$ yields a $\Lambda_{\mathfrak{p}}$-lattice $X_{\mathfrak{p}}$. Two $\Lambda$-lattices $X$ and $Y$ are said to be locally isomorphic, or in the same genus, if there is an isomorphism of $\Lambda_{\mathfrak{p}}$-lattices $X_{\mathfrak{p}}\cong Y_{\mathfrak{p}}$ for every $\mathfrak{p}$, in which case we write $X\vee Y$.
@@ 2.3. Locally free lattices and the principal genus
A $\Lambda$-lattice $X$ is said to be locally free of rank $n$ if $X\vee\Lambda^{\oplus n}$. Every locally free $\Lambda$-lattice is projective over $\Lambda$ by [CR81, (8.19)].
Define the principal genus $g(\Lambda)$ to be the class of locally free $\Lambda$-lattices of rank $1$. By [CR81, (31.7)], given $Y_{1},Y_{2},Y_{3}\in g(\Lambda)$, there exists $Y_{4}\in g(\Lambda)$ such that
(2.1) $Y_{1}\oplus Y_{2}\cong Y_{3}\oplus Y_{4}.$
Moreover, by [CR81, (31.14)], given $Y\vee\Lambda^{\oplus n}$, there exists $X\in g(\Lambda)$ such that
(2.2) $Y\cong\Lambda^{\oplus(n-1)}\oplus X.$
For $X\in g(\Lambda)$ we write $X]_{\mathrm{iso}}$ for the isomorphism class of $X$ and define
@@ 2.4. Stable isomorphism
Let $P(\Lambda)$ denote the class of finitely generated projective $\Lambda$-modules. For $X,Y\in P(\Lambda)$, we say that $X$ and $Y$ are stably isomorphic if
$X\oplus\Lambda^{\oplus m}\cong Y\oplus\Lambda^{\oplus m}\textrm{ for some }m\geq 0.$
A module $X\in P(\Lambda)$ is said to be stably free of rank $n$ if it is stably isomorphic to $\Lambda^{\oplus n}$. For $X\in P(\Lambda)$, let $X]_{\mathrm{st}}$ denote the stable isomorphism class of $X$. Note that by [CR81, (6.15)] and the discussion in §2.2, stably isomorphic $\Lambda$-lattices are locally isomorphic. In particular, stably free lattices are locally free.
@@ 2.5. Locally free class groups
By (2.1), given $X_{1},X_{2}\in g(\Lambda)$ there exists $X_{3}\in g(\Lambda)$ such that $X_{1}\oplus X_{2}\cong\Lambda\oplus X_{3}$. It is straightforward to check that the stable isomorphism class $X_{3}]_{\mathrm{st}}$ is uniquely determined by the classes $X_{1}]_{\mathrm{st}}$ and $X_{2}]_{\mathrm{st}}$. We define a binary operation on $\mathrm{Cl}(\Lambda)$ by
$X_{1}]_{\mathrm{st}}+[X_{2}]_{\mathrm{st}}=[X_{3}]_{\mathrm{st}}\textrm{ whenever }X_{1}\oplus X_{2}\cong\Lambda\oplus X_{3},$
where $X_{1},X_{2},X_{3}\in g(\Lambda)$. This is a well-defined commutative and associative operation, with identity element $\Lambda]_{\mathrm{st}}$. Moreover, inverses exist since by (2.1), given $X\in g(\Lambda)$ there exists $X^{\prime}\in g(\Lambda)$ such that $X\oplus X^{\prime}\cong\Lambda^{\oplus 2}$. Therefore $\mathrm{Cl}(\Lambda)$ is an abelian group. Moreover, $\mathrm{Cl}(\Lambda)$ is finite since $\mathrm{LF}_{1}(\Lambda)$ is finite and the canonical map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is surjective.